OpenStudy (anonymous):

Please help! Graph the circle with center at (3, -2), which also passes through the point (0, 2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (unanimoose):

Please don't tag everyone in your question.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

umm well have you ever used a program called geogebra because it might help because to tell you the truth I can help you on this sorry

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@Quan99

OpenStudy (cwrw238):

you can draw the graph manually by plotting the points (3,-2) the center and the point (0,2)

OpenStudy (cwrw238):

then using a compass place the point of the compass at the center and draw a circle through the other point.

OpenStudy (anonymous):

this might help I graphed it and I put 4 points on the circle

OpenStudy (cwrw238):

the distance between the 2 points is the radius of the circle and the equation of the circle will be (x -a )^2 + (y - b)^2 = r^2 where (a , b) is the center (3, -2) and r is the radius you have measured

OpenStudy (cwrw238):

@lexie<3 - the center is (3 , -2) not (-3,-2)

OpenStudy (anonymous):

oh soorrrryyyyy

OpenStudy (mathmath333):

use the diatance formula to find \(r\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{r=\sqrt{(3-0)^2+(-2-2)^2} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

would that be it? @cwrw238